OpenStudy (anonymous):

A circle has its center at (-2, -3) and a radius of 3 units. What is the equation of the circle? (x + 2)2 + (y + 3)2= 9 (x + 2)2 + (y - 3)2= 9 (x - 2)2 + (y - 3)2 = 3 (x + 2)2 + (y + 3)2 = 3

OpenStudy (ryan123345):

ok :)

OpenStudy (anonymous):

why whats wrong @ryan123345 ?

OpenStudy (stamp):

Equation of a circle with center (h, k) and radius r is\[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (ryan123345):

sorry i was just doing something ok not B or C

OpenStudy (stamp):

@irishchelsc So what is the equation if center (h, k) = (-2, -3) and your radius r = 3?

OpenStudy (anonymous):

isn't it (x-h)2 + (y-k)2

OpenStudy (anonymous):

h is the x coordinate and k is the y?

OpenStudy (ryan123345):

Yes :)

OpenStudy (ryan123345):

A

OpenStudy (anonymous):

Yay! thank god!

OpenStudy (anonymous):

that makes life so much easier :)

OpenStudy (ryan123345):

u didnt get my message?

OpenStudy (stamp):

If center (h, k) = (-2, -3) and radius r = 3, then substituting in the values provides:\[(x-(-2))^2+(y-(-3))^2=(3)^2\]Simplify to obtain:\[(x+2)^2+(y+3)^2=9\]

OpenStudy (anonymous):

thank you!! so much @ryan123345 and @stamp